the process is elimination by substitution. Equation(1) x+y over 2 minus x-y over 3 equals 8. Equation(2) x+y over 3 plus x-y over 4 equals 11
Answer by edjones(7311)

1)
(x+y)/2-(x-y)/3=8
6(x+y)/2-6(x-y)/3=6*8
3(x+y)-2(x-y)=48 No fractions.
3x+3y-2x+2y=48
x+5y=48
x=48-5y
.
2)
(x+y)/3+(x-y)/4=11
12(x+y)/3+12(x-y)/4=12*11
4(x+y)+3(x-y)=132
4x+4y+3x-3y=132
7x+y=132
7(48-5y)+y=132 substitute.
336-35y+y=132
336-336-34y=132-336
-34y=-204
-34y/-34=-204/-34
y=6
.
x=48-5y
x=48-5*6
=48-30
=18
.
Check:
(x+y)/3+(x-y)/4=11
(18+6)/3+(18-6)/4=11
8+3=11 true
.
Ed